# What Changed and What Could Have Changed? State-Change Counterfactuals for Procedure-Aware Video Representation Learning

### v2

@@ 2.1 Video Procedure-Aware Representations
Procedure understanding, which involves recognizing and localizing individual action steps and reasoning about their causal relationships, is a fundamental challenge in computer vision, and is important for applications such as step recognition [tang2019coin, damen2020epic, bansal2022my], temporal action segmentation [gtea, breakfast, sener2022assembly101, chen2025atars], and error detection [egoper]. Early work proposed learning procedure-aware video representations via exploiting the temporal consistency within videos [wang2015unsupervised, bansal2022my, xiao2022hierarchical, zhang2022unsupervised, qing2022learning] and temporal order [jenni2020video] as pre-text tasks. Some work incorporates task graphs as guidance to learn procedure-aware representations, either by constructing graphs based on knowledge bases [zhou2023procedure] or through learning [narasimhan2023learning, mavroudi2023learning, bansal2024united, seminara2024differentiable, ashutosh2024video]. Recently, several papers have attempted to align video features in datasets such as HowTo100M [miech2019howto100m] or Ego4D [grauman2022ego4d] with text descriptions extracted through automatic speech recognition [miech2020end, mavroudi2023learning], refined subtitles [lin2022learning] with the external database WikiHow [koupaee2018wikihow, mavroudi2023learning], or manually labeled annotations [kevin2022egovlp, zhong2023learning, ashutosh2023hiervl]. These text descriptions offer rich context of action steps and temporal relations, opening the door for procedure awareness learning. For example, Br-prompt [li2022bridge] generates text prompts from ground truth action labels to learn ordinal temporal features. PVRL [zhong2023learning] formulates future action steps as next-token prediction [devlin2019bert], enhancing temporal order between actions. HierVL [ashutosh2023hiervl] proposes hierarchical video-text alignment for both short clips and long videos.
@@ 3.3 Pretraining Objective:
$\mathcal{L}$ $=\frac{1}{|B|}\sum_{i\in B}\frac{-1}{|P(i)|}\sum_{p\in P(i)}\log\frac{\exp(f_{i}^{T}z_{p}/\tau)}{\sum_{n\in N(i)}\exp(f_{i}^{T}z_{n}/\tau)},$ (3)
@@ 4.5 Action Phase Classification & Retrieval Results
Table 4 shows that our model outperforms others on average across actions on both the ego+exo and ego settings, demonstrating its effectiveness in short-term procedures and strong view-point generalization. Interestingly, CLIP [radford2021learning] demonstrates strong performance. We conjecture this is due to its focus on short-term Full results with individual actions are available in Section LABEL:subsec:full in the supplementary material.
@@ 4.6 Ablation Study
Temporal Action Segmentation. One interesting insight on GTEA is that using only clip-level state changes and their counterfactuals (ID 3) achieves competitive performance on F1 and Edit metrics, which outperforms most other ablated models. Also, using only one type of video-level counterfactual (ID 4 and 5) may mislead the model to overfit a certain type of counterfactual, resulting in lower performance. By considering all the proposed states and counterfactuals together, we show that they are complementary to each other, thus achieving the best performance.
Error Detection. The results on EgoPER show that our proposed state changes and their counterfactuals can improve performance by combining them all together, verifying the effectiveness of recognizing erroneous labels. In addition, we observe interesting results showing that with only clip-level counterfactuals (ID 3), the model performs slightly worse than its baseline (ID 2). We assume this is because the model overfits short-term procedures and overlook long-term activities procedures and their counterfactual reasoning.

### v3

@@ 2.1 Video Procedure-Aware Representations
Procedure understanding, which involves recognizing and localizing individual action steps and reasoning about their causal relationships, is a fundamental challenge in computer vision, and is important for applications such as step recognition [tang2019coin, damen2020epic, bansal2022my], temporal action segmentation [gtea, breakfast, sener2022assembly101, chen2025atars], and error detection [egoper]. Early work proposed learning procedure-aware video representations via exploiting the temporal consistency within videos [wang2015unsupervised, bansal2022my, xiao2022hierarchical, zhang2022unsupervised, qing2022learning] and temporal order [jenni2020video] as pre-text tasks. Some work incorporates task graphs as guidance to learn procedure-aware representations, either by constructing graphs based on knowledge bases [zhou2023procedure] or through learning [narasimhan2023learning, mavroudi2023learning, bansal2024united, seminara2024differentiable, ashutosh2024video]. Recently, several papers have attempted to align video features in datasets such as HowTo100M [miech2019howto100m] or Ego4D [grauman2022ego4d] with text descriptions extracted through automatic speech recognition [miech2020end, mavroudi2023learning], refined subtitles [lin2022learning] with the external database WikiHow [koupaee2018wikihow, mavroudi2023learning], or manually labeled annotations [kevin2022egovlp, zhong2023learning, ashutosh2023hiervl]. These text descriptions offer rich context of action steps and temporal relations, opening the door for procedure-awareness learning. For example, Br-prompt [li2022bridge] generates text prompts from ground truth action labels to learn ordinal temporal features. PVRL [zhong2023learning] formulates future action steps as next-token prediction [devlin2019bert], enhancing temporal order between actions. HierVL [ashutosh2023hiervl] proposes hierarchical video-text alignment for both short clips and long videos.
@@ 3.3 Pretraining Objective:
$\mathcal{L}$ $=\frac{1}{|B|}\sum_{i\in B}\frac{-1}{|P(i)|}\sum_{z_{p}\in P(i)}\log\frac{\exp(f_{i}^{T}z_{p}/\tau)}{\sum_{z_{n}\in N(i)}\exp(f_{i}^{T}z_{n}/\tau)},$ (3)
@@ 4.5 Action Phase Classification & Retrieval Results
Table 4 shows that our model outperforms others on average across actions on both the ego+exo and ego settings, demonstrating its effectiveness in short-term procedures and strong view-point generalization. Interestingly, CLIP [radford2021learning] demonstrates strong performance. We conjecture this is due to the focus on short-term action in its image-to-text alignment. Full results with individual actions are available in Section LABEL:subsec:full in the supplementary material.
@@ 4.6 Ablation Study
Temporal Action Segmentation. One interesting insight on GTEA is that using only clip-level state changes and their counterfactuals (ID 3) achieves competitive performance, outperforming most other ablated models. Also, using only one type of video-level counterfactual (ID 4 and 5) may mislead the model to overfit a certain type of counterfactual, resulting in lower performance. By considering all the proposed states and counterfactuals together, we show that they complement each other, thus achieving the best performance.
Error Detection. The results on EgoPER show that our proposed state changes and their counterfactuals can improve performance by combining them all together, verifying the effectiveness of recognizing erroneous labels. In addition, we observe interesting results showing that with only clip-level counterfactuals (ID 3), the model performs slightly worse than its baseline (ID 2). We conjecture this is because the model overfits short-term procedures and overlook long-term activities procedures and their counterfactual reasoning.